With the aid of a specific example in each case, show what is meant by the
following types of spin coupling in 1 H NMR spectroscopy.
(i) geminal coupling,
(ii) vicinal coupling,
(iii) allylic coupling.
Answer #1

i) Geminal Coupling : The coupling of two non-equivalent protons on the same carbon atoms .These protons do not rotate around the other two carbon bonds as they are non-equivalent.(they have different types of atoms near them).It is called 2J coupling or two bond coupling as the protons are two bond away from each other

type:H-CH(R)R'' where R,R'' are different alkyl groups

example:H2C-(Cl)CH3

Here both H have different group near it so they are non-equivalent geminal protons

ii) vicinal coupling: The coupling of protons on adjacent or neighboring carbon atoms is called vicinal coupling .It is 3j coupling or three bond coupling as the protons are three bond away from each other

type:Ha-C-C-Hb

example:H-C(R)=C(R'')-H   R,R'' are different alkyl groups

iii)allylic coupling : The coupling between allylic protons or protons across one double bond (pi bond) and three single bonds is called four bond coupling or allylic coupling ,denoted by 4J

type:Ha-C=C-C-Hb (here Ha and Hb are allylic protons)

example: H-C=CH-CH2-H